OpenStudy (anonymous):

how do you convert -23pi/15 to degrees?

OpenStudy (travis):

pi = 180 degrees?

OpenStudy (anonymous):

yes

OpenStudy (nenadmatematika):

pi=180 degrees....so -23*180/15=-23*12=-276

OpenStudy (anonymous):

dude thank you so much !!!